OpenStudy (girlover):

Will fan and give medal. Write a specific formula to describe the variation: x varies inversely with the square of y; x = 6 when y = 4.

Nnesha (nnesha):

okay  this is formula for inversely

Nnesha (nnesha):

but your question is "x varies inversely with the SQUARE of y"

OpenStudy (girlover):

a. x= -24/sqroot y b. x=-24/y^2 c. x= 96/y^2 d. -96/sqroot y

OpenStudy (girlover):

@calculusxy

Nnesha (nnesha):

"x varies inversely with the SQUARE of y"  " now replace x and y by their values